A university bookstore recently sold a wirebound graph-paper notebook for $2.50, and a college-ruled notebook for $2.30. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $118.60. How many of each type were sold?
Answer by usyim88hk(158)

We can use Substitution to work on this problem:
~
Let x be the number of wirebound graph-paper notebook
Let y be the number of college-ruled notebook
~
So we now have an equation
x + y = 50
since the numbers of this two kinds of notebook add up to 50
~
now we can make another equation
2.5x + 2.3y = 118.6
~
2.5x means the earning of the sold wirebound graph-paper notebooks
2.3y means the earning of the sold college-ruled notebooks
and they add up to 118.6
~
Now we can solve for x
x + y = 50
x = 50 - y
~
Now substitute x into the other equation
2.5x + 2.3y = 118.6
2.5 (50 - y) + 2.3y = 118.6
(x is equal to 50 - y so the x in the second equation become 50-y)
~
Now solve for y
2.5 (50 - y) + 2.3y = 118.6
125 - 2.5y + 2.3y = 118.6
125 - 0.2y = 118.6
125 = 118.6 + 0.2y
6.4 = 0.2y
y = 32
So the total numebr of sold college-ruled notebooks is 32, while the total number of sold wirebound graph-paper notebooks is 50 - 32 = 18